OpenStudy (anonymous):

Let f(x)=4x+1and g(x)=x^2-2 (g*f)(-2)

OpenStudy (anonymous):

do i find the (g*f) (-2)?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

if it is, then here is the answer: (g*f)(x)=(4x+1)^2-2 (g*f)(x)=16x^2+8x-1 (g*f)(-2)=64-16-1 (g*f)(-2)=47

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no problem :D